[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Apr 18, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Apr 18 14:41:20 UTC 2017


HSU Shift Log for Tue Apr 18 2017 10:41:17 GMT-0400 (EDT)

Shift: H

Forecaster: L Avila

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues: P-Surge test was run for AL802017 at 1030 am EDT
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170418/792c6df5/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list